OpenStudy (anonymous):

Solving Systems By Substitution y=2x+1 y=x+3

jimthompson5910 (jim_thompson5910):

y=2x+1 x+3=2x+1 Now solve for x

OpenStudy (anonymous):

i need a little more help

jimthompson5910 (jim_thompson5910):

Let's solve for x x+3=2x+1 x-2x=1-3 -x = -2 x = -2/(-1) x = 2 Now use this to find y y=x+3 y=2+3 y = 5 So the solutions are x = 2 and y = 5

OpenStudy (anonymous):

thank you sooooo such

jimthompson5910 (jim_thompson5910):

yw